# Centre du Parc, Rue Marconi 19, CH-1920 Martigny, Switzerland

### v1

@@ 3.2 Tensor Train Decomposition
A $d$-th order tensor, denoted as $\bm{\mathcal{P}}\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$, serves as an extension of vectors (1-D) and matrices (2-D) to higher dimensions ($d$-D). Its configuration is defined by a tuple of integers $\bm{n}=(n_{1},\ldots,n_{d})$. Each element in the tensor, denoted as $\bm{\mathcal{P}}_{\bm{i}}$, is identified by $\bm{i}=(i_{1},\ldots,i_{d})$ as a set of $d$ integer variables representing the index in each dimension. The index set of the tensor is expressed as =$\mathcal{I}=\{{\bm{i}=(i_{1},\ldots,i_{d}):i_{k}\in\{1,\ldots,n_{k}\},k\in\{1,\ldots,d\}}\}$.
@@
Figure 4: TT decomposition extends matrix decomposition techniques to higher-dimensional arrays. In the TT format, accessing an element of a tensor involves multiplying the chosen slices (represented by green-colored matrices) of the core tensors (factors). The illustration provides examples for a 2nd order, 3rd order, and a 4th order tensor. This picture is adopted from Shetty et al. (2023)
@@ 3.3 Tensor-Train Distribution
This function approximation technique has been demonstrated within various optimization problems (Batsheva et al., 2023; Shetty et al., 2023). The problem of minimizing a cost function is first transformed into the task of maximizing an unnormalized probability density function $p(\cdot)$, represented as a tensor $\bm{\mathcal{P}}$ in TT format. Subsequently, a corresponding probability distribution, referred to as the TT distribution, is constructed as
@@
where $Z$ denotes the corresponding normalization constant. Leveraging the separable structure of the TT model allows for the efficient generation of exact samples from the TT distribution, eliminating the need to compute the normalization factor $Z$. Additional details about the sampling process can be found in the Appendix. Moreover, as illustrated in (Shetty et al., 2023), direct calculations of marginal and conditional distributions can be derived from the TT-cores.
@@ 4.1 Problem Definition
$\begin{gathered}\bm{U}^{*}_{t}=\underset{\hat{\bm{U}}_{t}}{\text{argmin}}\sum_{h=0}^{H-1}c_{h}(\hat{\bm{x}}_{t+h},\hat{\bm{u}}_{t+h})+c_{H}(\hat{\bm{x}}_{t+H}),\\ \text{s.t. }\bm{f}(\bm{x}_{t+h+1},\bm{u}_{t+h})\leq\bm{0},\quad\forall h\in\{0,1,\ldots,H-1\}.\end{gathered}$ (9)
Here, $\bm{x}_{t}$ denotes the system state at time step $t$, $\bm{f}(\cdot)\in\mathbb{R}^{w}$ represents the vector of $w$ constraints, and $c_{h}(\cdot)$ and $c_{H}(\cdot)$ denote the stage and terminal costs, respectively. For the sake of readability, task-dependent terms such as the goal position are not included in this equation. However, our method can be easily adapted to accommodate such cases. It is also important to note that the constraint set may incorporate prior knowledge that we expect the system to follow. For instance, in a pushing task, one constraint could ensure that actions taken at each step influence the object and prevent the robot from moving away from it beyond a certain threshold, as illustrated in Fig. 1-a.
@@ 4.2 Products of Experts
The first distribution $p^{o}$ captures the dynamic aspect of the task and ensures optimality (i.e., the optimality distribution, Expert 1), while the second distribution, $p^{f}$ represents the feasible area (i.e., the feasibility distribution, Expert 2).
Expert 1: We can solve (9) using the MPPI method by incorporating the hard constraint as a soft constraint within the cost function. This approach enables the system to gradually converge to the feasible area, although it may lead to time-consuming iterations and the rejection of numerous samples due to high-cost values. This expert addresses the following problem:
@@
where $c_{f}(\cdot)$ denotes the cost term associated with constraint violations in $\bm{f}$. The solution to (11) using the MPPI method can be formulated as
@@
Expert 2: To modify the sampling distribution of equation (12), we introduce another distribution $p^{f}(\cdot)$ that focuses on the feasible set :$\mathcal{F}:\{\bm{x},\bm{u}\,|\,\bm{f}(\bm{x},\bm{u})\leq 0\}$. We define a function $p$ to capture this feasible area as
$p(\bm{x}_{t},\bm{u}_{t})=\left\{\begin{array}[]{ll}1,&\bm{f}(\bm{x}_{t},\bm{u}_{t})\leq\bm{0}\\ 0,&\text{otherwise}\\ \end{array}\right.$ (13)
Various methods, including different generative models, can be employed to convert this function into a proper distribution. This article chooses the approach outlined in (Shetty et al., 2023). This method discretizes the function (13) and transforms it into an unnormalized distribution. The resulting distribution can be represented using tensor cores (as defined in Sec. 3.2), enabling a lower memory footprint and seamless integration with the other experts, along with a straightforward sampling method from the new distribution.
@@
$p^{f}(\bm{x}_{t},\bm{u}_{t})\rightarrow\bm{\mathcal{P}}_{f}(\hat{\bm{u}}_{t},\hat{\bm{x}}_{t})=\prod_{n=0}^{d_{x}}\bm{\mathcal{P}}^{n}_{{:,i_{\hat{x}_{n}},:}}\prod_{m=0}^{d_{u}}\bm{\mathcal{P}}^{d_{x}+m}_{{:,i_{\hat{u}_{m}},:}},$ (14)
where $\hat{x}_{n}$ and $\hat{u}_{m}$ are respectively $n$-th and $m$-th value of the state $\bm{x}_{t}$ and action $\bm{u}_{t}$ vectors. As discussed in (Shetty et al., 2023), it is possible to derive the conditional distribution in TT format $\bm{\mathcal{P}}_{f}(\hat{\bm{u}}_{t+h}|\hat{\bm{x}}_{t+h})$ from the joint distribution obtained in (14). Consequently, assuming that this distribution is time-independent (though it can be readily extended to a time-dependent version by introducing time as an additional variable into the system), we can define the second expert as
$p^{f}(\hat{\bm{U}}_{t}|\bm{x}_{t})=\prod_{h=0}^{H-1}\bm{\mathcal{P}}_{f}(\hat{\bm{u}}_{t+h}|\hat{\bm{x}}_{t+h}).$ (15)
@@ 4.3 TT-PoE-MPPI
$p(\hat{\bm{U}}_{t}|\bm{x}_{t})=\prod_{h=0}^{H-1}\mathcal{N}(\hat{\bm{u}}_{t+h}|\bm{\mu}_{t+h},\bm{\Sigma}_{t+h})\bm{\mathcal{P}}_{f}(\hat{\bm{u}}_{t+h}|\hat{\bm{x}}_{t+h}).$ (16)
@@
$p^{h}(\hat{\bm{u}}_{t+h}|\bm{x}_{t})=\mathcal{N}(\hat{\bm{u}}_{t+h}|\bm{\mu}_{t+h},\bm{\Sigma}_{t+h})\bm{\mathcal{P}}_{f}(\hat{\bm{u}}_{t+h}|\hat{\bm{x}}_{t+h}).$ (17)
@@
$p^{h}(\hat{\bm{u}}_{t+h}|\hat{\bm{x}}_{t+h})=\\ \mathcal{N}(\hat{\bm{u}}_{t+h}|\bm{\mu}_{t+h},\bm{\Sigma}_{t+h})\frac{\bm{\mathcal{P}}_{f}(\hat{\bm{u}}_{t+h},\hat{\bm{x}}_{t+h})}{Z},$ (18)
where $Z$ is a normalizing factor. This product can be computed efficiently by introducing a new tensor, $\bm{\mathcal{P}}_{\mathcal{N}}(\hat{\bm{u}}_{t+h})$, which represents the tensor form of the $\mathcal{N}(\cdot)$ function. Consequently, the product can be expressed as
$p^{h}(\hat{\bm{u}}_{t+h}|\hat{\bm{x}}_{t+h})=\frac{\bm{\mathcal{P}}_{\mathcal{N}}(\hat{\bm{u}}_{t+h})\bm{\mathcal{P}}_{f}(\hat{\bm{u}}_{t+h},\hat{\bm{x}}_{t+h})}{Z^{\prime}},$ (19)
where $Z^{\prime}$ is another normalizing factor. Although this product can be computed for any distribution, if we assume the covariance matrix is diagonal (i.e., independent actions in different dimensions), $\bm{\Sigma}_{t+h}=\text{diag}(\bm{\sigma}^{2}_{t+h})$, the computation becomes more efficient, allowing it to be performed directly at the core level. This assumption is common (Power and Berenson, 2023; Sacks and Boots, 2023) and is crucial as it enables efficient multiplication of these distributions.
@@
Figure 5: The pipeline employed in this paper integrates two distributions. Initially, the optimality distribution is transformed into TT-cores, followed by elementwise multiplication with the feasibility distribution cores corresponding to the action space. Subsequently, a new distribution is generated using the updated cores. In the illustration, white boxes represent elements equal to 1, while black boxes denote elements equal to 0. It is important to note that this visualization is simplified for clarity: in actual scenarios, the cores would not strictly be 0 or 1.
Algorithm 1 TT-PoE-MPPI
$\bm{x}_{t},\bm{\mu}_{t,\ldots,t+H-1},\bm{\sigma}_{t,\ldots,t+H-1}$
2:$\bm{u}^{*}_{t}$
3:$\hat{\bm{x}}_{t}\leftarrow\bm{x}_{t}$
4:for $h$ from 00 to $H-1$ do
5:     Calculate $\bm{\mathcal{P}}_{\mathcal{N}_{m}}$triangleright Eq. (21)
@@
7:     $\hat{\bm{u}}^{0}_{t+h},\ldots,\hat{\bm{u}}^{N}_{t+h}\leftarrow$Sample $N$ actions
@@
9:     $c^{0},\dots,c^{N}\leftarrow$Calculate cost values
10:     $\bm{\mu}_{t+h}\leftarrow$ update the mean value $\triangleright$ Eq. (4)
11:end for
12:$\bm{u}^{*}_{t}\leftarrow\bm{\mu}_{t}$
@@ 4.4 Decomposable feasible distribution
where $+$ and -- denote arithmetic operations between two TT models. These operations can further assist in solving more complex problems by focusing on simpler ones, allowing us to efficiently utilize the learned models instead of relearning them for each situation.
Figure 6: Examples of utilizing multiple experts to extract the feasible set. Left: Combining one expert that aligns the end-effector horizontally and another that positions the end-effector on a desired vertical line. The resulting distribution satisfies both constraints. The matrices show the corresponding configuration distribution with the darker values showing the Right: Similarly, combining models for obstacle avoidance and target reaching generates configurations (green lines) that meet both goals.
@@ 5 Results
We applied our methodology to several tasks, verifying the effectiveness of the proposed approach by conducting a comparative analysis with other baselines, namely MPPI, Proj-MPPI (when possible) and NFs-PoE-MPPI. In the MPPI baseline, constraints are integrated directly into the main cost function (i.e., as relaxed constraints), which aligns with the most widely accepted approach in the literature employing this method. In Proj-MPPI, each sample generated by the MPPI method was initially projected into the feasible region before being processed in the MPPI pipeline. The NFs-PoE-MPPI method leverages normalizing flows (NFs) (Dinh et al., 2016) to learn the feasibility distribution, taking inspiration from FlowMPPI (Power and Berenson, 2023). To learn this distribution with NFs, we used the same amount of data as for TT, ensuring that all data points were positive. We also maintained consistency in the number of data points across methods to enable a fair comparison. Details regarding the hyperparameters used in various learning methods across different tasks are provided in Table LABEL:tab:learning_data.

### v2

@@ 3.2 Tensor Train Decomposition
A $d$-th order tensor, denoted as $\bm{\mathcal{P}}\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$, serves as an extension of vectors (1-D) and matrices (2-D) to higher dimensions ($d$-D). Its configuration is defined by a tuple of integers $\bm{n}=(n_{1},\ldots,n_{d})$. Each element in the tensor, denoted as $\bm{\mathcal{P}}_{\bm{i}}$, is identified by $\bm{i}=(i_{1},\ldots,i_{d})$ as a set of $d$ integer variables representing the index in each dimension. The index set of the tensor is expressed as $\mathcal{I}=\{{\bm{i}=(i_{1},\ldots,i_{d}):i_{k}\in\{1,\ldots,n_{k}\},k\in\{1,\ldots,d\}}\}$.
@@
Figure 4: $T$ decomposition extends matrix decomposition techniques to higher-dimensional arrays. In the TT format, accessing an element of a tensor involves multiplying the chosen slices (represented by green-colored matrices) of the core tensors (factors). The illustration provides examples for a 2nd order, 3rd order, and a 4th order tensor. This picture is adopted from Shetty et al. (2024a) with permission from the authors.
@@ 3.3 Tensor-Train Distribution
The TT approximation technique has been previously applied to optimization problems (Batsheva et al., 2023; Shetty et al., 2024a). In this approach, the problem of minimizing a cost function is first transformed into the task of maximizing an unnormalized probability density function $p(\cdot)$, represented as a tensor $\bm{\mathcal{P}}$ in TT format. Subsequently, a corresponding probability distribution, referred to as the TT distribution, is constructed as
@@
where $Z$ denotes the corresponding normalization constant. Leveraging the separable structure of the TT model allows for the efficient generation of exact samples from the TT distribution, eliminating the need to compute the normalization factor $Z$. Additional details about the sampling process can be found in the Appendix. Moreover, as illustrated in (Shetty et al., 2024a), direct calculations of marginal and conditional distributions can be derived from the TT-cores.
@@ 4.1 Problem Definition
$\begin{gathered}\bm{U}^{*}_{t}=\underset{\hat{\bm{U}}_{t}}{\text{argmin}}\sum_{h=0}^{H-1}c_{h}(\hat{\bm{x}}_{t+h},\hat{\bm{u}}_{t+h},{\bm{\theta}})+c_{H}(\hat{\bm{x}}_{t+H},{\bm{\theta}}),\\ \text{s.t. }\bm{f}(\hat{\bm{x}}_{t+h+1},\hat{\bm{u}}_{t+h},{\bm{\theta}})\leq\bm{0},\quad\forall h\in\{0,1,\ldots,H-1\}.\end{gathered}$ (9)
Here, $\bm{x}_{t}$ denotes the system state at time step $t$, $\bm{f}(\cdot)\in\mathbb{R}^{w}$ represents the vector of $w$ constraints, and $c_{h}(\cdot)$ and $c_{H}(\cdot)$ denote the stage and terminal costs, respectively. $\bm{\theta}$ denotes task-dependent terms such as obstacle locations, target positions, etc. For clarity, we omit these terms in the remainder of the paper; however, they are still incorporated in our experiments, as detailed in the Appendix. It is important to note that the constraint set may incorporate prior knowledge that we expect the system to follow. For instance, in a pushing task, one constraint could ensure that actions taken at each step influence the object and prevent the robot from moving away from it beyond a certain threshold, as illustrated in Fig. 1-a.
@@ 4.2 Products of Experts
The first distribution, $p^{o}$, captures the dynamic aspect of the task and ensures optimality (i.e., the optimality distribution, Expert 1), while the second distribution, $p^{f}$, represents the feasible area (i.e., the feasibility distribution, Expert 2).
Expert 1: We can solve (9) using the MPPI method by incorporating the constraint within the cost function. This approach enables the system to gradually converge to the feasible area, although it may lead to time-consuming iterations and the rejection of numerous samples due to high-cost values. This expert addresses the following problem:
@@
where $c_{f}(\cdot)$ denotes the cost term associated with constraint violations in $\bm{f}$. Although Expert 1 is primarily responsible for modeling the optimality distribution, the inclusion of the additional cost term $c_{f}(\cdot)$ offers several advantages. First, by assigning significantly higher values to this term relative to others, infeasible samples are effectively assigned zero weight via (5), thereby removing them from consideration. Secondly, it allows us to express the objective in a more general form, consistent with how similar formulations are structured in related literature, thereby making Expert 1 directly comparable to prior approaches. Third, this added term provides flexibility to account for certain constraints that may emerge during task execution but are not explicitly captured by Expert 2 (described later). Finally, from a practical standpoint, incorporating $c_{f}(\cdot)$ into Expert 1 can enhance efficiency by guiding the optimization using relevant constraint information, even when a separate feasibility expert is also present. In this paper, we follow prior works (Jankowski et al., 2023; Bhardwaj et al., 2022; Power and Berenson, 2023) by assigning a large weight to this term in order to more effectively reject infeasible solutions. Bhardwaj et al. (2021) provide a detailed analysis of how increasing this penalty improves feasibility under tight constraints.
The solution to (11) using the MPPI method can be formulated as
@@
Expert 2: Handling constraints exclusively through a rejection mechanism, for example by assigning high values to the $c_{f}$ term, is generally inefficient. It wastes computational resources by generating and then discarding infeasible trajectories. Our proposed method addresses this by introducing a feasibility expert that proactively biases the sampling distribution toward feasible regions. This improves the quality of samples before they are evaluated and potentially rejected, leading to several key benefits as discussed in this paper.
To modify the sampling distribution, we introduce another distribution $p^{f}(\cdot)$ that focuses on the feasible set $\mathcal{F}:\{(\hat{\bm{x}},\hat{\bm{u}}\,|\,\bm{f}(\bm{x},\bm{u})\leq 0\}$. We define a function $p_{\text{feas}}$ to capture this feasible area as
$p^{f}(\hat{\bm{x}}_{t},\hat{\bm{u}}_{t})\propto p_{\text{feas}}(\hat{\bm{x}}_{t},\hat{\bm{u}}_{t})=\exp\Big(-\frac{\text{ReLU}\Big(\bm{f}(\hat{\bm{x}}_{t},\hat{\bm{u}}_{t})\Big)}{\lambda}\Big),$ (13)
where $\lambda$ is a temperature parameter. Various methods, including different generative models, can be employed to convert this function into a proper distribution. This article chooses the approach outlined in (Shetty et al., 2024a). This method discretizes the function (13) and transforms it into an unnormalized distribution. The resulting distribution can be represented using tensor cores (as defined in Sec. 3.2), enabling a lower memory footprint and seamless integration with the other experts, along with a straightforward sampling method from the new distribution.
@@
$p^{f}(\hat{\bm{x}}_{t},\hat{\bm{u}}_{t})\rightarrow\bm{\mathcal{P}}_{f}(\hat{\bm{u}}_{t},\hat{\bm{x}}_{t})=\prod_{n=0}^{d_{x}}\bm{\mathcal{P}}^{n}_{{:,i_{\hat{x}_{n}},:}}\prod_{m=0}^{d_{u}}\bm{\mathcal{P}}^{d_{x}+m}_{{:,i_{\hat{u}_{m}},:}}$, (14)
where $\hat{x}_{n}$ and $\hat{u}_{m}$ are respectively $n$-th and $m$-th value of the state $\hat{\bm{x}}_{t}$ and action $\hat{\bm{u}}_{t}$ vectors. As discussed in (Shetty et al., 2024a), it is possible to derive the conditional distribution in TT format $\bm{\mathcal{P}}_{f}(\hat{\bm{u}}_{t+h}|\hat{\bm{x}}_{t+h})$ from the joint distribution obtained in (14). Consequently, assuming that this distribution is time-independent (though it can be readily extended to a time-dependent version by introducing time as an additional variable into the system), we can define the second expert as
$p^{f}(\hat{\bm{U}}_{t}|\bm{x}_{t})\propto\prod_{h=0}^{H-1}\bm{\mathcal{P}}_{f}(\hat{\bm{u}}_{t+h}|\hat{\bm{x}}_{t+h}).$ (15)
@@ 4.3 TT-PoE-MPPI
$p(\hat{\bm{U}}_{t}|\bm{x}_{t})\propto\prod_{h=0}^{H-1}\mathcal{N}(\hat{\bm{u}}_{t+h}|\bm{\mu}_{t+h},\bm{\Sigma}_{t+h})\bm{\mathcal{P}}_{f}(\hat{\bm{u}}_{t+h}|\hat{\bm{x}}_{t+h}).$ (16)
@@
$p^{h}(\hat{\bm{u}}_{t+h}|\bm{x}_{t})\propto\mathcal{N}(\hat{\bm{u}}_{t+h}|\bm{\mu}_{t+h},\bm{\Sigma}_{t+h})\bm{\mathcal{P}}_{f}(\hat{\bm{u}}_{t+h}|\hat{\bm{x}}_{t+h}).$ (17)
@@
$p^{h}(\hat{\bm{u}}_{t+h}|\hat{\bm{x}}_{t+h})\propto\\ \mathcal{N}(\hat{\bm{u}}_{t+h}|\bm{\mu}_{t+h},\bm{\Sigma}_{t+h})\bm{\mathcal{P}}_{f}(\hat{\bm{u}}_{t+h},\hat{\bm{x}}_{t+h}).$ (18)
This product can be computed efficiently by introducing a new tensor, $\bm{\mathcal{P}}_{\mathcal{N}}(\hat{\bm{u}}_{t+h})$, which represents the tensor form of the $\mathcal{N}(\cdot)$ function. Consequently, the product can be expressed as
$p^{h}(\hat{\bm{u}}_{t+h}|\hat{\bm{x}}_{t+h})\propto\bm{\mathcal{P}}_{\mathcal{N}}(\hat{\bm{u}}_{t+h})\bm{\mathcal{P}}_{f}(\hat{\bm{u}}_{t+h},\hat{\bm{x}}_{t+h}).$ (19)
Although this product can be computed for any distribution, if we assume the covariance matrix is diagonal (i.e., independent actions in different dimensions), $\bm{\Sigma}_{t+h}=\text{diag}(\bm{\sigma}^{2}_{t+h})$, the computation becomes more efficient, allowing it to be performed directly at the core level. This assumption is common (Power and Berenson, 2023; Sacks and Boots, 2023) and is crucial as it enables efficient multiplication of these distributions.
@@
Figure 5: The pipeline employed in this paper integrates two distributions. Red, and blue colors indicate the cores related to the state and action variables, respectively. Initially, the optimality distribution is transformed into $T$-cores, followed by elementwise multiplication with the feasibility distribution cores corresponding to the action space. Subsequently, a new distribution is generated using the updated cores. In the illustration, white boxes represent elements equal to 1, while darker boxes denote elements equal to 0. It is important to note that this visualization is simplified for clarity: in actual scenarios, the cores would not strictly be 0 or 1.
@@
Algorithm 1 $T$-PoE-MPPI
1:Require: $\bm{x}_{t},\bm{\mu}_{t,\ldots,t+H-1},\bm{\sigma}_{t,\ldots,t+H-1}$
2: Return: $\bm{u}^{*}_{t}$
[…]
4:for $h$ from 0 to H1H-1 do
5:  Calculate $\bm{\mathcal{P}}_{\mathcal{N}_{m}}$ $\triangleright$ Eq. (21)
@@
7:  $\hat{\bm{u}}^{0}_{t+h},\ldots,\hat{\bm{u}}^{N}_{t+h}\leftarrowSample$ NN actions
@@
9:end for
[…]
@@ 4.4 Decomposable feasible distribution
As a side note, when the feasible regions of two constraints overlap, we can use a relaxed combination via summation rather than multiplication as
$\bm{\mathcal{P}}^{f}=\bm{\mathcal{P}}^{f}_{1}+\bm{\mathcal{P}}^{f}_{2},$ (23)
where ++ denotes arithmetic operations between two TT models. While a product-of-experts formulation is often preferred, the additive combination retains the same high-probability overlap of feasible regions and thus remains effective at guiding the sampler toward the feasible set. Moreover, this relaxed sum-of-experts requires fewer ranks and can be more practical in high-dimensional settings.
@@
Figure 6: Examples of utilizing multiple experts to extract the feasible set. Left: Combining one expert that aligns the end-effector horizontally and another that positions the end-effector on a desired vertical line. The resulting distribution satisfies both constraints. The matrices show the corresponding configuration distribution with the darker values showing the desired configurations. Right: Similarly, combining models for obstacle avoidance and target reaching generates configurations (green lines) that meet both goals.
@@
These operations can further assist in solving more complex problems by focusing on simpler ones, allowing us to efficiently utilize the learned models instead of relearning them for each situation.
@@ 5 Results
We applied our methodology to several tasks, verifying the effectiveness of the proposed approach by conducting a comparative analysis with other baselines, namely MPPI, Proj-MPPI (when possible) and NFs-PoE-MPPI. In the MPPI baseline, constraints are integrated directly into the main cost function (i.e., as relaxed constraints), which aligns with the most widely accepted approach in the literature employing this method. In Proj-MPPI, each sample generated by the MPPI method was initially projected into the feasible region before being processed in the MPPI pipeline. The NFs-PoE-MPPI method leverages normalizing flows (NFs) (Dinh et al., 2016) to learn the feasibility distribution, taking inspiration from FlowMPPI (Power and Berenson, 2023). To learn this distribution with NFs, we used the same amount of data as for TT, ensuring that all data points were positive. We maintained consistency in the number of data points across methods to enable a fair comparison. Details regarding the hyperparameters used in various learning methods across different tasks are provided in Table LABEL:tab:learning_data.
The core problem, particularly the optimality distribution, is identical for all baselines. Specifically, each method employs a high-penalty term with the same value within the cost function to reject infeasible samples. This mechanism is critical for the constrained tasks considered in this paper; without it, even a few infeasible samples that score well on the unconstrained cost could degrade performance. This approach is consistent with prior work (Jankowski et al., 2023; Bhardwaj et al., 2022; Power and Berenson, 2023). When comparing baselines, we therefore focus on evaluating the unique components of each method, such as the feasibility expert in PoE-MPPI, rather than the shared rejection mechanism. This separation allows us to independently assess the contribution of these additional components.
@@
Finally, we demonstrate that our framework scales to high-dimensional problems and can seamlessly integrate with alternative optimality distributions, such as the diffusion-based DIAL-MPC (Xue et al., 2025).